# Euclid Quick Data Release (Q1) point-like X-ray selected sources

### v1

@@ 5 Characterisation and classification of CTPs
Once the CTPs have been identified, it is essential to categorise the sources in order to study the underlying physical processes and populations. The primary distinction is between extragalactic sources (such as galaxies or faint and bright AGN) and galactic sources (including stars and compact objects). In the following section, we outline the methodology used for classifying the sources and the validation tests conducted to ensure the accuracy of the classification, the results of which are summarised in Footnote 8.
@@ 5.1 Galactic and extragalactic sources
The four Euclid bands alone lack the discriminatory power required to reliably separate Galactic from extragalactic sources, given the broadness of the optical filter, where most of the differences within the SED occur. As demonstrated in Euclid Collaboration: Bisigello et al. (2024), no combination of Euclid-only photometry achieves satisfactory fractions of purity and completeness for this task. Moreover, the ground-based photometry added to the Q1 catalogues is not completely cross-calibrated at this stage, and differences in quality and bands available differ from field to field. For this reason, we decided to use the photometry available in LS10, providing us with well-calibrated and reasonable depth in from DECam Legacy Survey observations, including data from the Dark Energy Survey (DES, Dark Energy Survey Collaboration et al. 2016), Beijing-Arizona Sky Survey (BASS, Zou et al. 2017), and the Mayall -band Legacy Survey (MzLS, Silva et al. 2016), as well as W[1,2,3,4] from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE, Mainzer et al. 2011; Lang 2014; Meisner et al. 2017). We match the 11 286 CTPs to LS10 using a positional tolerance and find an association for 68% of our sources, see Footnote 8. Next, we apply a cut of S/N ¿ 3 in the , , , and W1 bands to minimise the relative impact of photometric uncertainties. The selected sources are then plotted on the W1 versus colour-colour plane, to distinguish between Galactic and extragalactic objects (see Fig. 9). Using the separation line defined in Salvato et al. (2022), we preliminarily classify sources based on their location relative to the threshold. Objects above the threshold are flagged as extragalactic, while those below are classified as Galactic.
@@ 5.2 Probabilistic classifier approach
To overcome this limitation, we follow the methodology outlined in Sect. 4.1.1, training an RF classifier using the same Euclid-only features listed in Footnote 6. However, this time we label the sample based on the classification derived from the versus plot, setting Galactic objects (396) to ‘1‘ and extragalactic objects (1000) to ‘0‘. Using the trained model, we assign a Galactic or extragalactic label to each Euclid CTP, providing a probabilistic classification for robust downstream analyses.
We present the respective confusion matrix in Fig. 10 and note a recall of 92%, with a limited fall-out of 4%. For more details on the feature correlations see Appendix A. To further validate the RF performance, we examine the galactic probability distribution () for the unique CTP candidates sample. Less than 10% of sources have , indicating effective separation of Galactic and extragalactic populations. A comparison with the Euclid Collaboration: Matamoro Zatarain et al. (2025) classification reveals that 97% of sources classified as Galactic there, also have , confirming good agreement. Additionally, the colour-coding of Fig. 11 reveals that most sources classified as being Galactic make up the bright end of the sample, with the most intense colour gradient surrounding and a couple of sources towards high interestingly appearing increasingly fainter again, probably due to different populations of stars, subject to their location in the extragalactic sky.
@@ 6 Redshifts
The catalogues of CTPs we make available in this paper include spectroscopic redshifts (spec-s), where available, by matching their coordinates to a compilation of publicly available redshifts (Kluge et al. 2024). However, this compilation is rich in duplications, hence the details of the cleaning are described in Sect. 3.1 of Saxena et al. (2024). While spec-s provide highly accurate distance measurements, they are inherently limited to a subset of typically brighter objects. As a result only 16% of the CTP sample benefits from publicly available spec- coverage (refer to Footnote 8).
@@ 6.1 Photometric redshifts
To achieve a more complete characterization of the CTPs, photometric redshifts (photo-s) are essential. However, achieving reliable photo-s for AGN remains significantly more challenging than for inactive galaxies, despite notable advances in recent years (see Salvato et al. 2019, for a review).
@@
The primary difficulty lies in the overwhelming emission produced by the accretion onto the SMBH, which often masks the spectral features of the host galaxy, crucial for breaking colour-redshift degeneracies. The current version of photo-s released for Q1 assumes the sources to be inactive galaxies and does not account for the possible contribution from an active nucleus. The absence of AGN SEDs in the template library makes it difficult to accurately constrain their contributions using photometry alone, resulting in miscalculations of the physical parameters for each source derived indirectly. Instead, studies have demonstrated that incorporating aperture or pixel-based analyses can significantly enhance the accuracy of such photo- estimates. Consequently, to compute reliable photo-s for AGN-dominated objects in a dedicated effort, we utilise PICZL (Roster et al. 2024), an ML algorithm that significantly improves photo- estimation for AGN by directly predicting redshifts from imaging data, eliminating the need for manual feature extraction or additional surveys. Building on its predecessor, Circlez (Saxena et al. 2024), which analysed aperture-based photometric variations, PICZL advances to pixel-level resolution, leveraging the raw images for improved precision. As a result, for sources bright enough to successfully be matched to the LS10 catalogue, we are able to produce reliable photo-s (refer to Footnote 8). This is possible since the training sample constructed for PICZL is formed mostly by X-ray selected AGN detected in LS10, enabling us to use the same model. As in the original work, the uncertainty of the photo-s in terms of 1 error, increases with the faintness of the sources (see Roster et al. 2024, for more details).
@@
For all EDF CTPs of an X-ray source detected in LS10, PICZL provides full redshift posterior distributions (PDZ) derived from de-reddened calibrated images in the , , , and bands, supplemented by , , , and catalogue-based aperture photometry without imposing any S/N cuts. The method is agnostic to the morphological classification of the source, applying equally to extended and point-like objects. As output, PICZL provides the dominant mode of the PDZ as a point prediction, along with uncertainties quantified as 1 and 3 confidence intervals. Notably, we compute photo-s also for those CTPs matched to LS10, which we believe to be Galactic. This approach accounts for the inherent uncertainty in the probabilities produced by our model and classification threshold. To ensure flexibility for users, we include both classifications and photo- estimates in the released catalogue (see Sect. 8). This allows users to tailor their selection criteria to their preferences while retaining access to the full set of photo- values.
@@ 6.2 Photo-z quality
To evaluate the reliability of our photo- estimates, we employ standard metrics commonly implemented in the literature, including (a) the accuracy as a measure of the scatter between the prediction and truth values of the sample, defined as
and (b) the fraction of outliers, , which quantifies the proportion of sources whose redshift estimates deviates significantly from their corresponding spectroscopic redshift. Specifically, this is defined as the fraction of objects for which
We evaluated the performance of the photo-s by comparing them to the spec-s, as shown in Fig. 12. For the subsample of sources with ¡ 21.5, which corresponds to the brighter and more reliable sources in our data set, we find that the fraction of outliers is approximately 5%, with a of roughly 0.04 as opposed to and observed for the entire sample. Additionally, the catastrophic outliers at are faint, galaxy-dominated objects ( ¿ 24) with large photometric errors in LS10. In scientific studies involving photometric redshifts, it is crucial to account for such uncertainties and in turn, photometric redshift algorithms should therefore make sure to provide reliable errors. PICZL achieves this effectively as discussed in detail in Roster et al. (2024). In the right panel of Fig. 12 this is evidenced by the smaller errors closely tracing the identity line, while larger errors appropriately account for sources of uncertainty.
In Fig. 13, we present the cumulative distribution of CTP band magnitudes along with the respective fractions of sources with spectroscopic and/or and photometric redshifts. The figure reveals that the sample is almost complete in redshift up to around . For fainter sources, photo- estimation is either not possible due to the limited number of photometric data points and significant errors, or the source, though clearly detected in Euclid, is not visible in LS10 due to its shallower depth. In the future, when we have upgraded PICZL to be applied to Euclid images and have integrated data from the Legacy Survey of Space and Time (LSST; Ivezic et al. 2019), photometric redshifts for AGN in Euclid will dramatically improve.
@@ 7 Properties of the CTP sample
Figure 14 illustrates the magnitude distribution of our CTPs, grouped by the match between EDFs and the corresponding X-ray survey (4XMM, CSC2, and eROSITA DR1, respectively). The 4XMM and CSC 2.0 distributions show their ability to detect fainter X-ray sources, which correlates with the identification of fainter optical CTPs. In contrast, the eROSITA DR1 distributions exhibit brighter median magnitudes, consistent with the shallowness of the survey (see Fig. 2). Additionally, the figure suggests a possible bimodal distribution at the faint end in the optical band for both 4XMM and CSC 2.0, at least in EDF-N and EDF-S. The majority of ources in the second, faint peak of the distribution show X-ray fluxes , with no obvious drop off in p_any, and the distribution will be studied further in other papers.
@@ 7.1 Colour-redshift relation
Redshifts, spectroscopic and photometric, are available for 79% of the CTPs, allowing us to analyse their Euclid colours as function of redshift, as shown in Fig. 15. The colour-coded kernel density is plotted along the y-axis dimension within each bin, providing a visual trace of how the physical colours of the CTP sample vary as a function of redshift.
A notable feature in the data is the detection of the Balmer break at , particularly prominent in colours including the filter. Unsurprisingly, NIR-only colour combinations display relatively few discernible features (compare e.g., figure 7 in Temple et al. 2021). As iterated before, the Euclid filters alone lack the fidelity to catch variations in the SED of sources, a caveat that we will be able to overcome in combination with upcoming surveys, especially LSST.
@@ 7.2 X-ray luminosities
For the rest of the analysis, we restrict the CTP sample to those 7345 extragalactic sources with a soft X-ray flux S/N 2 in order to reduce potential spurious X-ray detection. In Fig. 16, we observe a clear distinction between pointed observations taken with XMM-Newton and Chandra at various depths versus the homogenous all-sky coverage provided by eROSITA. The complete coverage of eROSITA results in a tighter correlation between luminosity and redshift, despite the presence of a few points away from the general trend, which have low S/N (thus indicating potential spurious sources). Another reason for their offset from the general trend could be an incorrect photometric redshift. On the other hand, XMM-Newton and Chandra catalogues include a mixture of pointed observations resulting in a less smooth distribution between redshift and luminosity.
@@
In all three panels, we indicate with open circle sources with low S/Nr, potentially corresponding to those highlighted red in the left panel of Fig. 12, representing examples of increased redshift uncertainties. Notably, most of the results shown, focus on objects with redshifts of . Sources with higher redshifts, while present, should be approached with caution due to potential inaccuracies. However, the future availability of LSST photometry will enable the exploration of this parameter regime by probing fainter sources with smaller errors, thus improving the reliability of redshift estimates for these populations.
In studies such as that of Georgakakis & Nandra (2011), AGN are typically defined as sources with X-ray luminosities exceeding , since normal galaxies may be an important source of contamination below this limit (Lehmer et al. 2012; Aird et al. 2015). However, it is important to note that this threshold was established based on pencil-beam surveys, suggesting that it should be revisited now that wide-area surveys and deep optical data are accessible. X-ray luminosities of AGN in the soft X-ray band (0.5 – 2 keV) can vary significantly, depending on the AGN type and redshift. In literature, values can range from or even less when considering low-luminosity AGN (Hasinger et al. 2005; Ebrero et al. 2009; Civano et al. 2012; Kawamuro et al. 2013). Applying the criterion of , we classify 6243 out of the 6295 sources with computed X-ray luminosities as AGN.
@@
A more inclusive approach is presented in Georgakakis et al. (2007) for example, where X-ray detected starburst galaxy candidates, though more likely in case of Chandra as opposed to eROSITA (Kyritsis et al. 2025), can represent about 20% of the X-ray source population in the luminosity interval . This fraction is also likely to be an upper limit because some of the galaxy candidates may turn out to be low–luminosity AGN, such that distinguishing between these sources can be challenging. Using our sources from XMM-Newton, where both soft and hard (2 – 12 keV)999We assume that at zero order, the luminosities between (2 – 10) and (2 – 12) keV are approximately the same. band fluxes are available, we show that, to retain all sources in this parent catalogue that are classified as AGN using the hard band, a soft-band luminosity threshold of is required (refer to Fig. 17). Given the results of both all-sky and pencil beam surveys, at soft luminosities in the range , AGN still appear to represent the dominant population. Nonetheless, future studies are required to determine the nature of these low-luminosity AGN at soft energies. Applying this more inclusive threshold defined in Fig. 17 to our extragalactic CTP sample, 6292 of the 6295 sources that we were able to compute X-ray luminosities for, meet the AGN definition criterion.

### v2

@@ 5 Characterisation and classification of CTPs
Once the CTPs have been identified, it is essential to categorise the sources in order to study the underlying physical processes and populations. The primary distinction is between extragalactic sources (such as galaxies or faint and bright AGN) and galactic sources (including stars and compact objects). In the following section, we outline the methodology used for classifying the sources and the validation tests conducted to ensure the accuracy of the classification, the results of which are summarised in Footnote 9.
@@ 5.1 Galactic and extragalactic sources
The four Euclid bands alone lack the discriminatory power required to reliably separate Galactic from extragalactic sources, given the broadness of the optical filter, where most of the differences within the SED occur. As demonstrated in Euclid Collaboration: Bisigello et al. (2024), no combination of Euclid-only photometry achieves satisfactory fractions of purity and completeness for this task. Moreover, the ground-based photometry added to the Q1 catalogues is not completely cross-calibrated at this stage, and differences in quality and bands available differ from field to field. For this reason, we decided to use the photometry available in LS10, providing us with well-calibrated and reasonable depth in from DECam Legacy Survey observations, including data from the Dark Energy Survey (DES, Dark Energy Survey Collaboration et al. 2016), Beijing-Arizona Sky Survey (BASS, Zou et al. 2017), and the Mayall -band Legacy Survey (MzLS, Silva et al. 2016), as well as W[1,2,3,4] from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE, Mainzer et al. 2011; Lang 2014; Meisner et al. 2017). We match the 11 286 CTPs to LS10 using a positional tolerance and find an association for 82% of our sources, see Footnote 9. Next, we apply a cut of S/N ¿ 3 in the , , , and W1 bands to minimise the relative impact of photometric uncertainties. The selected sources are then plotted on the W1 versus colour-colour plane, to distinguish between Galactic and extragalactic objects (see Fig. 9). Using the separation line defined in Salvato et al. (2022), we preliminarily classify sources based on their location relative to the threshold. Objects above the threshold are flagged as extragalactic, while those below are classified as Galactic.
@@ 5.2 Probabilistic classifier approach
To overcome this limitation, we followed the methodology outlined in Sect. 4.1.1, training an RF classifier using the same Euclid-only features listed in Footnote 7. However, this time we labelled the sample based on the classification derived from the versus plot, setting Galactic objects (396) to ‘1‘ and extragalactic objects (1000) to ‘0‘. Using the trained model, we assigned a Galactic or extragalactic label to each Euclid CTP, providing a probabilistic classification, , for robust downstream analyses.
We present the respective confusion matrix in Fig. 10 and note a recall of 92%, with a limited fall-out of 4%. For more details on the feature correlations, see Appendix A. To further validate the RF performance, we examine the galactic probability distribution () for the unique CTP candidates sample. Less than 10% of sources have , indicating effective separation of Galactic and extragalactic populations. A comparison with the Euclid Collaboration: Matamoro Zatarain et al. (2025) classification reveals that 97% of sources classified as Galactic there, also have , confirming good agreement. Additionally, the colour-coding of Fig. 11 reveals that most sources classified as being Galactic make up the bright end of the sample, with the most intense colour gradient surrounding and a couple of sources towards high interestingly appearing increasingly fainter again, probably due to different populations of stars, subject to their location in the extragalactic sky.
@@ 6 Redshifts
The catalogues of CTPs we make available in this paper include spectroscopic redshifts (spec-s), where available. This was done by matching their coordinates to a compilation of publicly available redshifts (Kluge et al. 2024, Igo et al. in prep.). However, this compilation is rich in duplications; hence, the details of the cleaning are described in Sect. 3.1 of Saxena et al. (2024). While spec-s provide highly accurate distance measurements, they are inherently limited to a subset of typically brighter objects. As a result, only 16% of the CTP sample benefits from publicly available spec- coverage (refer to Footnote 9).
@@ 6.1 Photometric redshifts
To achieve a more complete characterisation of the CTPs, photometric redshifts (photo-) are essential. However, achieving reliable photo-s for AGN remains significantly more challenging than for inactive galaxies, despite notable advances in recent years (see Salvato et al. 2019, for a review). The current Q1 photo-s (Euclid Collaboration: Tucci et al. 2025), derived using Phosphoros (Paltani et al. in prep.), are estimated from total fluxes and fine-tuned for inactive galaxies. This approach neglects the photometric signatures of active nuclei, whose emission can dominate the observed flux and hide key host galaxy features, crucial for breaking colour-redshift degeneracies (Salvato et al. 2019). Instead, studies have demonstrated that incorporating aperture or pixel-based analyses can significantly enhance the accuracy of such photo- estimates. Consequently, to compute reliable photo-s for AGN-dominated objects in a dedicated effort, we utilise PICZL (Roster et al. 2024), an ML algorithm that significantly improves photo- estimation for AGN by directly predicting redshifts from imaging, eliminating the need for manual feature extraction or combination of various surveys. Building on its predecessor, Circlez (Saxena et al. 2024), which analysed aperture-based photometric variations, PICZL advances to pixel-level resolution, leveraging the raw images for improved precision. As a result, for sources bright enough to successfully match the LS10 catalogue, we are able to produce reliable photo-s (refer to Footnote 9). As in the original work, the uncertainty of the photo-s in terms of 1 error increases with the faintness of the sources (see Fig. 12 and Roster et al. 2024, for more details).
This study does not incorporate Euclid photometry in the photo- estimation, as doing so would require a dedicated pipeline, including re-training on Euclid imaging and assembling a spectroscopic AGN training set from the Q1 field. Given the small and incomplete nature of such a sample, especially at the faint end, this effort is deferred until DR1 becomes available. An additional challenge lies in the heterogeneous photometric coverage across Q1: while Euclid provides uniform measurements in , , , and , these are complemented by ground-based data of varying depth and filter sets, leading to inconsistencies in photo- quality. Consequently, to ensure robustness and comparability, we opted to rely on established methods, with PICZL offering high completeness and reliable redshifts for 82% of the X-ray sources detected in LS10101010Future PICZL versions will make use Euclid imaging, possibly combined with LSST (Ivezic et al. 2019). (see Fig. 13). Nonetheless, we recognise the potential of Euclid photometry to enhance our performance, particularly for faint sources and at high redshifts (Graham et al. 2020). As illustrated in the left panel of Fig. 12, numerous faint sources are dispersed horizontally around , as no strong spectral breaks can be captured by the Legacy filters around this redshift (Roster et al. 2024). Access to Euclid ’s NIR bands , , and will enable more accurate redshift recovery in this range.
@@
For all EDF CTPs of an X-ray source detected in LS10, PICZL provides full redshift posterior distributions (PDZ), derived from de-reddened calibrated images in the , , , and bands, supplemented by , , , and catalogue-based aperture photometry without imposing any S/N cuts. The method is agnostic to the morphological classification of the source, applying equally to extended and point-like objects. As output, PICZL provides the dominant mode of the PDZ as a point prediction, along with uncertainties quantified as 1 confidence intervals. Notably, we compute photo-s also for those CTPs matched to LS10, which we believe to be Galactic. This approach accounts for the inherent uncertainty in the probabilities produced by our model and classification threshold. To ensure flexibility for users, we include both classifications and photo- estimates in the released catalogue (see Sect. 8). This allows users to tailor their selection criteria to their preferences while retaining access to the full set of photo- values.
@@ 6.2 Photo-z quality
To evaluate the reliability of our photo- estimates, we employed standard metrics commonly implemented in the literature, including (a) the accuracy as a measure of the scatter between the prediction and truth values of the sample, which is defined as
and (b) the fraction of outliers, , that quantifies the proportion of sources whose redshift estimates deviate significantly from their corresponding spectroscopic redshift. Specifically, this is defined as the fraction of objects for which
We evaluated the performance of the photo-s by comparing them to the spec-s, as shown in the middle panel of Fig. 12. For the subsample of bright sources with ¡ 21.5, corresponding to those that have smaller photometric errors within our dataset, we find a fraction of outliers of 5.4%, with a of roughly 0.04 as opposed to and observed for the entire sample. The catastrophic outliers at appear as faint ( ¿ 24), galaxy-dominated objects with larger photometric errors in LS10 and highly degenerate PDZs. In scientific studies involving photo-s, it is crucial to account for such uncertainties and in turn, photo- algorithms should provide reliable error estimates. PICZL achieves this effectively as shown in the right panel of Fig. 12, with small errors closely following the identity line and larger ones capturing photometric uncertainties. Despite occasional higher uncertainties, point estimates remain robust, ensuring that for applications such as luminosity functions, the entire PDZs remain valuable.
In Fig. 13, we present the cumulative distribution of CTP band magnitudes along with the corresponding fractions of sources for which spec- and/or and photo-s from PICZL are available. The figure demonstrates that redshift completeness is very high up to around , beyond which the availability of reliable photo- estimates begins to decline. In addition to spectroscopic selection effects, this is due to fainter sources which, while clearly detected in Euclid, often lack sufficient optical coverage or exhibit photometric uncertainties too large to yield robust photo- values. Future releases of PICZL, adapted to directly process Euclid imaging and incorporating the deeper and more complete photometry expected from the Legacy Survey of Space and Time (LSST; Ivezic et al. 2019), will markedly enhance photo- performance for fainter AGN.
We compared our photo- estimates with those from Duncan (2022), who combined ML and SED fitting to derive redshifts for all extragalactic sources in the 8th data release of the Legacy Survey (LS8). In EDFN, where LS10 photometry corresponds roughly to that of LS8, PICZL achieves a slight improvement in (from 0.10 to 0.08) and a notable reduction in (from 25% to 14%), without applying magnitude cuts. In the two additional fields covered by LS10-South (Zenteno et al. 2025), where PICZL also benefits from deeper WISE and added i-band photometry, the improvements are even more pronounced, with improving from 0.09 to 0.05, and dropping from 23% to 12%. Likewise, while the photo-s from Zhou et al. (2021, 2023), computed with LS10 photometry, perform extremely well for inactive galaxies, a trend similar to Duncan (2022) is observed for AGN, where the accuracy is known to decline111111with reaching up to 33% in some subsets (priv. comm.). For high-redshift ( ¿ 3) X-ray selected AGN, we compared our estimates to those of Pouliasis et al. (2025), computed using LS10 supplemented with near-infrared photometry from the VISTA Hemisphere Survey (VHS, McMahon et al. 2013). Of the 45 sources in common, only two have spec- for which both photo- agree, preventing any meaningful additional comparison. Instead, we can assess our photo- quality for X-ray selected AGN at the faint end by benchmarking them against studies that employ significantly deeper and broader photometric datasets.
@@ 7 Properties of the CTP sample
Building on this approach, we note that a similar comparison has already been presented in Roster et al. (2024), where PICZL photo-s were evaluated against the XMM-SERVS catalogues from Chen et al. (2018) and Ni et al. (2021). Here, we extend this strategy by comparing our photo- results with the 7 Ms Source Catalog from the Deep Field-South (CDFS, Luo et al. 2017). This catalogue is particularly well-suited for faint-end validation, as it includes photo-s derived using high-quality multi-band data, largely informed by Hsu et al. (2014). Utilising up to 38 intermediate and broad bands spanning the ultraviolet (GALEX, Martin et al. 2005) to the mid-infrared (Spitzer/IRAC, Cardamone et al. 2010), their work has resulted in one of the most comprehensive SED-fitting datasets for faint X-ray-selected sources to date. To ensure a proper one-to-one comparison between sources, we match our CTP catalogue to that of Luo et al. (2017) using the respective CTP coordinates. We recover the same CTPs across the full sample, with 97.5% of sources showing offsets below 08, with the largest observed separation being 18. Indicating robust matches, this consistency highlights the overall quality of our CTPs that can be attributed to the effectiveness of the Euclid-based prior used with NWAY, providing a solid foundation for future applications. For the 400 sources with spec-s matched between this work and Luo et al. (2017), we compare the two datasets by evaluating and as a function of -band magnitude from the Wide Field Imager (WFI; Giavalisco et al. 2004). As depicted in Fig. 23, PICZL achieves a lower fraction of outliers than Luo et al. (2017) for 22.3, though at larger due to using only four shallower optical and WISE mid-IR bands. However, even up to PICZL extrapolation results in outlier fractions of , which is lower compared to, for example, eROSITA / eFEDS (Salvato et al. 2022), where the fraction is %, despite using many more and deeper bands. Given the overall quality of Luo et al. (2017), we decided to extend redshift completeness in this work by incorporating 378 unique photo-s from the 7 Ms Source Catalog.
Figure 14 illustrates the magnitude distribution of our CTPs, grouped by the match between EDFs and the corresponding X-ray survey (4XMM DR14, CSC 2.0, and eROSITA DR1, respectively). The 4XMM and CSC2 distributions show their ability to detect fainter X-ray sources, which correlates with the identification of fainter optical CTPs. In contrast, the eROSITA DR1 distributions exhibit brighter median magnitudes, consistent with the shallowness of the survey (see Fig. 2). Additionally, the figure suggests a possible bimodal distribution at the faint end in the optical band for both 4XMM and CSC 2.0, at least in EDF-N and EDF-S. The majority of sources in the second, faint peak of the distribution show X-ray fluxes , with no obvious drop off in p_any, and the distribution will be studied further in other papers.
@@ 7.1 Colour-redshift relation
Redshifts, spectroscopic and photometric, are available for 81% of the CTPs, allowing us to analyse their Euclid colours as a function of redshift, as shown in Fig. 15. The colour-coded kernel density is plotted along the y-axis dimension within each bin, providing a visual trace of how the physical colours of the CTP sample vary as a function of redshift.
A notable feature in the data is the detection of the Balmer break at , particularly prominent in colours including the filter. Unsurprisingly, NIR-only colour combinations display relatively few discernible features (compare e.g., figure 7 in Temple et al. 2021). As iterated before, Euclid filters alone lack the fidelity to catch variations in the SED of sources, a caveat that we will be able to overcome in combination with upcoming surveys, especially LSST.
@@ 7.2 X-ray luminosities
For the rest of the analysis, we restrict the CTP sample to those 7345 extragalactic sources with a soft X-ray flux S/Nx 2 in order to reduce potential spurious X-ray detections. In Fig. 16, we observe a clear distinction between pointed observations taken with XMM-Newton and Chandra at various depths versus the homogenous all-sky coverage provided by eROSITA. The complete coverage of eROSITA results in a tighter correlation between luminosity and redshift, despite the presence of a few points away from the general trend, with S/N. Another reason for their offset from the general trend could be erroneous photo-. On the other hand, XMM-Newton and Chandra catalogues include a mixture of pointed observations, resulting in a less smooth distribution between redshift and luminosity.
@@
In all three panels, we indicate with open circles sources with low S/Nr, presumably representing examples of increased redshift uncertainties. Notably, most of the results shown focus on objects with redshifts of . Sources with higher redshifts, while present, should be approached with caution due to potential inaccuracies. However, the future availability of LSST photometry will enable the exploration of this parameter regime by probing fainter sources with smaller errors, thus improving the reliability of redshift estimates for these populations.
@@
In studies such as that of Georgakakis & Nandra (2011), AGN are typically defined as sources with X-ray luminosities exceeding , since normal galaxies may be an important source of contamination below this limit (Lehmer et al. 2012; Aird et al. 2015). However, it is important to note that this threshold was established based on pencil-beam surveys, suggesting that it should be revisited now that wide-area surveys and deep optical data are accessible. X-ray luminosities of AGN in the soft X-ray band (0.5–2 keV) can vary significantly, depending on the AGN type and redshift. In literature, values can range from or even less when considering low-luminosity AGN (Hasinger et al. 2005; Ebrero et al. 2009; Civano et al. 2012; Kawamuro et al. 2013). According to Georgakakis et al. (2007), for example, X-ray detected starburst galaxy candidates, though more likely in case of Chandra as opposed to eROSITA (Kyritsis et al. 2025), can represent about 20% of the X-ray source population in the luminosity interval . This fraction is also likely to be an upper limit because some of the galaxy candidates may turn out to be low-luminosity AGN, such that distinguishing between these sources can be challenging. Consequently, with sources from XMM-Newton, where both soft and hard (2–12 keV)121212We assume that at zero order, the luminosities between (2–10) and (2–12) keV are approximately the same. band fluxes are available, we show that, to retain all sources in this parent catalogue that are classified as AGN using the hard band limit, a soft-band luminosity threshold of is required (refer to Fig. 17). Given the results of both all-sky and pencil beam surveys, at soft luminosities in the range , AGN still appear to represent the dominant population. Nonetheless, future studies are required to determine the nature of these low-luminosity AGN at soft energies. Applying this more inclusive threshold defined in Fig. 17 to our extragalactic CTP sample, all of the 6653 (91%) sources that we were able to compute X-ray luminosities for, meet the AGN definition criterion, 49 of which have .